# Large-Scale Training of Graph Neural Networks for Optimal Markov-Chain Partitioning Using the Kemeny Constant

### v2

@@ 4 Results & Discussion
However, examples derived from numerical simulations of bio-molecular systems often result in noisy derived MCs with dense connections, such as in the pentapeptide case. Here, as shown in Fig. 7 and Table 2, the simplest GraphSAGE and linear decoder method outperforms all other methods similarly to the 1D example above. The PVTC method and PCCA+ also cannot correctly identify the optimal clustering. In this example our method finds the optimal clustering less frequently than in the synthetic examples of SBM networks. This could be due to the increased number of low $\Delta K$ partitionings available, which provide many local minima, which are more difficult to escape. MCMC methods struggle with this same issue, having difficulty finding the relevant areas of partitioning space. The dense nature of the clusters means, it is difficult for PCCA+-based methods to identify the most optimal clustering do not differ much in terms of meta-stability. Nevertheless, the optimal PCCA+ clustering is remarkably close to the optimal Kemeny constant-based clustering. The larger models suffer from a vanishing gradient due to the heavily conditional nature of optimizing all node partition probabilities at the same time, resulting in an inability to find the optimal clustering even once while nevertheless providing a better result on average than the GraphSAGE and LD model. Still, they are outclassed by the PCCA+, and all but the GraphSAGE with transformer decoder model provide worse performance than PVTC, highlighting the efficiency of simpler networks for this kind of task. While further analysis of learning rate adjustments could provide an effective solution to train larger and more complex models on this problem, it is clear that training simple networks is both quicker and more reliable to find the optimal clustering. Feature comparisons on the GraphSAGE + LD example show the success of the method in finding the optimal partitioning over 1000 individual runs with all features tested except for using the eigenvectors (Fig. 4b). MFPTs show the best performance here, finding the optimal solution the greatest number of times. Most likely due to the computational simplicity in finding effective mappings from the MFPTs to optimized partitionings, as the value of Eq. (20) directly depends upon them. The poor performance of the eigenvectors is most likely down to numerical stability, as some of them are very small 64-bit floats.
@@
One advantage MCMC methods such as the PVTC have over our method is the ability to be initialised from any arbitrary partitioning. The results shown are for the PVTC when given no initial clustering information, however, when the method is initialised from the PCCA+ solution it consistently finds the optimal solution in all occasions. To try to replicate this behaviour, we changed our initialisation function $\mathcal{I}$ to pretrain the models to replicate the PCCA+ solution using mean-squared error loss that is minimal for the optimal PCCA+ solution. Once the GNN is trained to this specific loss function and identifies the PCCA+ clustering, it is then further trained to the Kemeny loss of Eq. (20). This pretraining, however, did not improve performance of the GNN optimisation (data not shown).
@@
Figure 6: (a) Comparison of the PCCA+ solution with the optimal $\Delta K$ clustering for the pentapeptide system. Networks are displayed using a force-directed layout weighted with the MFPTs. Nodes which differ between the two clusters are highlighted with a yellow border.
@@
(b) Violin plot distributions for 1000 experiments with each set of nine features. The corresponding distribution of final $\Delta K$ values is shown as a violin plot, along with a blue dashed line for the optimal $\Delta K$ value.
@@
Figure 7: The violin plot distribution of $\Delta K$ values found from training 1000 copies of each neural network architecture on the pentapeptide system. The blue dashed line indicates the value associated with the optimal clustering.
@@
We also performed a hyperparameter search on our GNN network to identify the optimal depth and network sizes for the training. The parameter search was performed on the pentapeptide example. It highlights the problem that the network’s effectiveness is highly dependent upon the chosen hyper-parameters of the neural network and the size of the network. The best parameter space for a given problem must be sufficiently large to be able to express the wide multitude of possible $\mathbf{S}$ matrices, while also being small enough to allow for effective optimisation in a feasible time. This is highlighted by the poorer performance of most networks featuring four linear layers in the decoder, where most of the time the method either fails to find the optimal solution or does so rarely it is classified as an outlier. Future work can further verify whether a varying or learnable adjustment to the learning rate could provide a reasonable improvement, as this would allow networks to more efficiently probe the relevant regions of parameter space. Alternatively, further work could evaluate if it is possible to quantify the relationship between the size of the graph being clustered and the most optimal network size.
Figure 8: Parameter grid search on the pentapeptide example. 1000 networks were initialised for each set of parameters, and the resulting $\Delta K$ distribution is shown. The grey background indicates the parameters identified as optimal.
@@ 5 Conclusion
We describe a GNN-based network clustering optimisation method, which uses the computing infrastructure and tooling that has risen around GNNs to provide a working basis for difficult optimization problems on graphs. The benefits are numerous, as GNNs provide an ample Ansatz for beginning optimization for partitioning phenomena in situations where otherwise finding an efficient representation for the task proves difficult.
@@
We provide a novel implementation of a GP approach to cluster graphs based on kinetic optimization criteria. Here we use the Kemeny constant to obtain metastable kinetically optimal clusters. We outline a general methodology for training GNNs to optimize graph partitionings and compare the performance of four different architectures across three different kinetic network types, using several different node features. Our results show that smaller and simpler networks often show superior performance in our test cases, as larger networks become more difficult to optimize, and often cannot find the optimal solutions. Testing the method on SBM graphs with clear clusters, we can attain similar results to existing methods, such as PCCA+. Frequently GNN optimizations even outperform the PCCA+ solution in terms of $\Delta K$ on SBM graphs with ill-defined clusters. However, SBM networks are generally easier to cluster compared to networks defined from kinetic processes. A linear chain corresponding to an analytic 1D free energy profile also poses a difficult task for GNNs due to the low connectivity, while PCCA+ in this case can produce the optimal solution. We show how we can adapt the node features the GNN uses and the number of runs to alleviate this problem, and improve the performance. However, the optimal clustering is only found in about 5% of the attempts in only one GNN architecture, and therefore improving the reliability of the method still requires future work. The outline of our general methodology, introducing the strategy defining generator $\mathcal{I}$, culling $\mathcal{C}$, and termination $\mathcal{T}$ functions, provides an ample framework for future work in this direction. Our MD-simulation derived MSM for a pentapeptide has an optimal $\Delta K$ clustering that seems to better capture metastable states than the PCCA+-derived clusters. Only one of the four architectures can find this optimal solution in less than 3% of the time, highlighting the problem of GNNs, which further work on $\mathcal{I}$ function can hopefully provide a solution for. Additionally, GNNs do not provide a straightforward way to initialize the search from a close-to-optimal solution, which makes the PVTC algorithm much more efficient.
In summary, we have shown the validity of using GNNs to effectively cluster MCs and described a general framework which is straightforward to generalize to both different cluster optimization criteria and graph-like structures. We have proposed a GNN-based methodology for GP that can be further generalized and explored in future work. It remains to be seen how to implement more complex parameter space sampling strategies to aid the identification of more optimal solutions, taking inspiration from similar techniques used for enhanced sampling in MD.

### v3

@@ 4 Results & Discussion
In this example, our method finds the optimal clustering less frequently than in the synthetic examples of SBM networks. This could be due to the increased number of low $\Delta K$ partitionings available, which provide many local minima which are more difficult to escape. However, this is not a problem unique to the GNN method, as the PTVC method encounters the same issue, having difficulty finding the relevant areas of partitioning space and is unable to find the optimal clustering in this example using random initialization. The dense and interconnected nature of the nodes means it is difficult for PCCA+-based methods to identify the most optimal clustering as well, as more optimal clusters may not differ much in terms of meta-stability. The larger models suffer from a vanishing gradient due to the heavily conditional nature of optimizing all node partition probabilities at the same time, resulting in an inability to find the optimal clustering even once while nevertheless providing a better result on average than the GraphSAGE and LD model. The best clusterings found by the GraphSAGE and TD models are shown in Fig. S4 for reference. Still, they are outclassed by the PCCA+, and all but the GraphSAGE with linear decoder model provide worse performance than PTVC, highlighting the efficiency of simpler networks for this kind of task.
While further analysis of learning rate adjustments could provide an effective solution to train larger and more complex models on this problem, it is clear that training simple networks is both quicker and more reliable in finding the optimal clustering. Feature comparisons on the GraphSAGE + LD example show the success of the method in finding the optimal partitioning over 1000 individual runs, with all features tested except for using the eigenvectors (Fig. 6b). MFPTs show the best performance here, finding the optimal solution the greatest number of times. Most likely due to the computational simplicity in finding effective mappings from the MFPTs to optimized partitionings, as the value of Eq. (23) directly depends upon them. The poor performance of the eigenvectors is most likely down to numerical stability, as some of them are very small 64-bit floats.
@@
One advantage of the PTVC method, which uses an MCMC-based algorithm, over our method, is the ability to be initialized from any arbitrary partitioning. The results shown are for the PTVC when given no initial clustering information, however, when the method is initialized from the PCCA+ solution it consistently finds the optimal solution on all occasions. In the GNN models, we also changed our initialization function $\mathcal{I}$ by pretraining the models to replicate the PCCA+ solution using mean-squared error loss that is minimal for the optimal PCCA+ solution. Once the GNN is trained to this specific loss function and identifies the PCCA+ clustering, it is then further trained using the Kemeny loss of Eq. (23). This pretraining, however, did not improve the performance of the GNN optimization (data not shown).
@@
Figure 6: (a) Comparison of the PCCA+ solution with the optimal $\Delta K$ clustering for the pentapeptide system, which has DBI scores of 0.823 and 0.758 respectively. Networks are displayed using a force-directed layout weighted with the MFPTs. Nodes which differ between the two clusters are highlighted with a yellow border.
@@
(b) Violin plot distributions for 1000 experiments with each set of nine features. The corresponding distribution of final $\Delta K/KC$ values is shown as a violin plot, along with a blue dashed line for the optimal $\Delta K/KC$ value.
@@
Figure 7: (a) A bar-plot showing the DBI values associated with the optimal $\Delta K$ partitionings found by each method used. (b) A violin distribution plot of $\Delta K/K$ values acquired from training 1000 copies of each neural network architecture on the pentapeptide system. The blue dashed line indicates the value associated with the optimal clustering.
Fig. 8 shows the results of a hyperparameter search using the best-performing simple GNN model to identify the optimal parameters for the network on the pentapeptide example. It highlights the problem that the network’s effectiveness is highly dependent upon the chosen hyper-parameters of the neural network and the size of the network. The best parameter space for a given problem must be sufficiently large to be able to express the wide multitude of possible $\mathbf{S}$ matrices, while also being small enough to allow for effective optimization in a feasible time. This is highlighted by the poorer performance of most networks featuring four linear layers in the decoder, where most of the time the method either fails to find the optimal solution or does so rarely it is classified as an outlier. Future work can further verify whether a varying or learnable adjustment to the learning rate could provide a reasonable improvement, as this would allow networks to more efficiently probe the relevant regions of parameter space. Alternatively, further work could evaluate if it is possible to quantify the relationship between the size of the graph being clustered and the most optimal network size.
@@
PCCA+ 258.86 ( $\pm$ 0.0) 258.86 N/A
Figure 8: Parameter grid search on the pentapeptide example. 1000 networks were initialized for each set of parameters, and the resulting $\Delta K/KC$ distribution is shown. The grey background indicates the parameters identified as optimal.
@@ 5 Conclusion
We describe a GNN-based network clustering optimization method, which uses the computing infrastructure and tooling that has risen around GNNs to provide a working basis for difficult optimization problems on graphs. The benefits are numerous, as GNNs provide an ample Ansatz for beginning optimization for partitioning phenomena in situations where otherwise finding an efficient representation for the task proves difficult.
We provide a novel implementation of a GP approach to cluster graphs based on kinetic optimization criteria. Here we use the KC to obtain metastable kinetically optimal clusters. We outline a general methodology for training GNNs to optimize graph partitionings and compare the performance of four different architectures across three different kinetic network types, using several different node features. Our results show that smaller and simpler networks often show superior performance in our test cases, as larger networks become more difficult to optimize, and often cannot find the optimal solutions. However, as we are only interested in the best solution with the maximal Kemeny constant, multiple evaluations can lead to the correct global solution, even if only obtained less frequently in individual initializations. GraphSAGE layers combined with a linear decoder achieved both greater consistency and accuracy than any tested attention-based alternatives.
@@
To show that our proposed method provides a viable alternative to existing methods in terms of the coarse grained systems, we compare optimal clusterings obtained with those identified using the PCCA+ clustering. It is important to note, however, that as PCCA+ uses a different kinetics-based objective function, the task of identifying the optimal clustering using the KC-based measure is independent of this comparison. Testing the method on SBM graphs generated to have a clear optimal partitioning, we can attain similar results to PCCA+, but with more optimal values of the Kemeny constant. A linear chain corresponding to an analytic 1D free energy profile also poses a difficult task for GNNs due to the low connectivity, while PCCA+ in this case can produce the optimal solution. We show how we can adapt the node features the GNN uses and the number of runs to alleviate this problem, and improve the performance. However, the optimal clustering is only found in about 5% of the attempts in only one GNN architecture, and therefore improving the reliability of the method still requires future work. The outline of our general methodology, introducing the strategy defining generator $\mathcal{I}$, culling $\mathcal{C}$, and termination $\mathcal{T}$ functions, provides an ample framework for future work in this direction.
Our MD-simulation derived MSM for a pentapeptide is the most complex example we considered here. It has a slightly more optimal $\Delta K$ clustering that seems to better capture metastable states than the PCCA+-derived clusters, as also seen by comparing the corresponding DBI scores. Only one of the four architectures can find this optimal solution in less than 3% of the time, highlighting the problem of GNNs, which further work on $\mathcal{I}$ functions can hopefully provide a solution for. Additionally, GNNs do not provide a straightforward way to initialize the search from a close-to-optimal solution, which makes the PTVC algorithm more efficient.
In this work, we explored a number of node features that were used by the GNNs to enable efficient optimization. The MFPT-based features were some of the most successful ones, as these directly relate to the definition of $\Delta K$. However, for examples derived from biomolecular simulations, features that are specifically relevant to physical and structural observables (angles, distances, energy terms, etc.) could outperform the general features used here. This is a promising avenue for future directions.
In summary, we have shown the validity of using GNNs as an effective method to solve the GP problem applied to MCs, alongside describing a general framework outlining the strategy for optimizing partitionings using GNNs in parallel on graph-like structures. We propose using such a framework to solve the many different and difficult non-convex optimization tasks which are commonplace in computational fields using graph structures, taking full advantage of the speed, availability, and efficiency of modern ML tooling and associated hardware. The general training paradigm outlined provides a working basis for generalizing this strategy to different problems, and the results highlight the usefulness of simpler GNN architectures when solving such optimization problems. It remains to be seen how this can be improved by implementing more complex parameter space sampling strategies to aid the identification of more optimal solutions, taking inspiration from similar techniques used for enhanced sampling in MD.